OpenStudy (anonymous):

Use division to determine whether the ratios form a proportion. 12 : 44 and 3 : 11 A. No. They do not form a proportion. B. Yes. They form a proportion.

OpenStudy (anonymous):

@jabez177

OpenStudy (anonymous):

Is It B?

OpenStudy (cheyennecoffelt):

12/44=0.27272727... 3/11=0.27272727...

OpenStudy (cheyennecoffelt):

That is a proportion

OpenStudy (anonymous):

Thank You c:

OpenStudy (cheyennecoffelt):

or multiply each by a factor they share. 3*4=12 and 11*4=44, which is 12:44